Bed & Bath, a retailing company, has two departments—Hardware and Linens. The company’s most recent monthly contribution format income statement follows:

Department
Total Hardware Linens
Sales $ 4,290,000 $ 3,110,000 $ 1,180,000
Variable expenses 1,390,000 971,000 419,000
Contribution margin 2,900,000 2,139,000 761,000
Fixed expenses 2,160,000 1,320,000 840,000
Net operating income (loss) $ 740,000 $ 819,000 $ (79,000 )

A study indicates that $372,000 of the fixed expenses being charged to Linens are sunk costs or allocated costs that will continue even if the Linens Department is dropped. In addition, the elimination of the Linens Department will result in a 12% decrease in the sales of the Hardware Department.

Required:

What is the financial advantage (disadvantage) of discontinuing the Linens Department?
Answer #1

Contribution margin for Hardware would be=2,139,000*(1-0.12)=$1882320

Less:Fixed cost for Hardware=(1,320,000)

Net operating income for Hardware=$562320

Less:Unavoidable fixed cost for Linens=(372,000)

Total net operating income would be=$190320

Hence financial disadvantage=740,000-190320

=$549680